i have a parallelogram with side lengths 9cm and 3 cm. how do i find the height?..so then i can find the area.: i have a parallelogram with side lengths 9cm and 3 cm. how do i find the height?..so then i can find the area.
Answer by Earlsdon(3816)

Since you provide no more information than the lengths of the sides of the parallelogram, we have to assume that the parallelogram in your problem is a rectangle, in which case, the area is given by: A = L*W where L = 9cm and W = 3cm.
So the area is: A = 9*3 = 27 sq.cm.